# Pulse Design of Baseband Flux Control for Adiabatic Controlled-Phase Gates in Superconducting Circuits

### v1

@@ Abstract
Despite progress towards achieving low error rates with superconducting qubits, error-prone two-qubit gates remain a bottleneck for realizing large-scale quantum computers. Therefore, a systematic framework to design high-fidelity gates becomes imperative. One type of two-qubit gate in superconducting qubits is the controlled-phase (CPHASE) gate, which utilizes a conditional interaction between higher energy levels of the qubits controlled by a baseband flux pulse on one of the qubits or a tunable coupler. In this work, we study an adiabatic implementation of CPHASE gates and formulate the design of the control trajectory for the gate as a pulse-design problem. We show in simulation that the Chebyshev-based trajectory can, in certain cases, enable gates with leakage error lower by an average of roughly $6\%$ when compared to the widely used Slepian-based trajectory.
@@ II.4.2 Constraint on duration
$P_{e}=\frac{\displaystyle\bigg{|}\int\tilde{g}(\tau)e^{-i\Delta\tau}\textrm{d}\tau\bigg{|}^{2}}{4}=\frac{\displaystyle\big{|}G(i\Delta)\big{|}^{2}}{4}\ .$ (7)
@@ IV Comparison between Chebyshev- and Slepian-based trajectories
We have formulated the CPHASE gate design problem into a pulse design problem and further transformed it into the design of a control trajectory $\tilde{g}[n]$. Special attention is paid to the comparisons between the Chebyshev-based trajectories and the Slepian-based trajectories that will be defined in this section.
@@
We propose the Chebyshev-based trajectory $\tilde{g}_{\textrm{ch2}}[n]$ as an alternative to the Slepian-based trajectory $\tilde{g}_{\textrm{sl2}}[n]$. The major characteristic difference of the two control trajectories is that $\tilde{g}_{\textrm{ch2}}[n]$ has equiripple sidelobe amplitude for all sidelobes, while $\tilde{g}_{\textrm{sl2}}[n]$ has decreasing sidelobe amplitude. The argument is that we can allow higher sidelobe amplitudes in larger frequency components as long as they stay below a specified threshold, and therefore we can in turn decrease the concentration in smaller frequency components. In this way, $\tilde{g}_{\textrm{ch2}}[n]$ can be designed to obtain a smaller $\omega^{*}$ while maintaining a sidelobe amplitude below some threshold. Here, $\omega^{*}$ denotes the smallest attainable frequency under certain constraints given some leakage threshold, as we specify in Statement 3. This eventually leads to the fact that the Chebyshev-based trajectory can be designed to be shorter than the Slepian-based trajectory.
@@ V.1 Setting and procedure
In our simulation, we choose the parameters $\omega_{2}=4.7$ GHz, $\alpha_{1}=\alpha_{2}=-300$ MHz, and $g=14.142$ MHz. We set $\omega_{1}=5.8$ GHz initially and tune $\omega_{1}$ to perform a CZ gate. These parameters form a typical parameter set for transmon qubits and the operation of a CZ gate. In order to perform a CZ gate, we detune $\omega_{1}$ so that $\omega_{1}+\omega_{2}\approx 2\omega_{1}+\alpha_{1}$. If we move exactly to the degeneracy point of diabatic states |$\ket{11}$ and |$\ket{20}$, then we have $\omega_{1}=5.0$ GHz. We vary $\omega_{1}$ from $5.8$ GHz to approximately $5.0$ GHz depending on the amplitude of the control pulse, and then back to $5.8$ GHz.
@@ V.2.1 A comparison example
Potential operating points are considered to be the insensitive points to pulse duration, i.e., the “sweet spots” of the leakage error lobes as shown by the green squares and purple dots in Fig. 3. In addition, if there is a slight deviation in pulse duration at those points, we would be almost surely have a lower leakage error regardless of the direction of the deviation. We determine the best operating point to be the one with the smallest pulse duration, indicated by the red circle in Fig. 3. Since $\tilde{g}_{\textrm{ch2}}[n]$ generally pushes the leakage error lower in the range of relatively smaller pulse duration, we are able to achieve an operating point with lower leakage error while also maintaining a similar or even smaller pulse duration compared to $\tilde{g}_{\textrm{sl2}}[n]$. The operating points for $\tilde{g}_{\textrm{sl2}}[n]$ and $\tilde{g}_{\textrm{ch2}}[n]$ are $t_{d}=47.0$ ns with $P_{e}=10^{-4.66}$ and $t_{d}=46.1$ ns with $P_{e}=10^{-4.72}$, respectively.
@@ V.2.2 An aggregate of comparison examples
We observe similar results to that in Section V.2.1 for a majority of the comparison examples. We find that $\tilde{g}_{\textrm{ch2}}[n]$ can be designed to achieve a lower leakage error than its counterpart $\tilde{g}_{\textrm{sl2}}[n]$ by roughly $6\%$ in certain cases, while also maintaining smaller pulse duration by an average of $0.6$ ns. In several abnormal cases, where there appears an unusually small leakage lobe before the first main leakage lobe, $\tilde{g}_{\textrm{sl2}}[n]$ in fact induces lower leakage error with smaller pulse duration. This indicates that simulation verification is important once a control trajectory is designed because there can be discrepancies between analysis and simulation due to some approximation.
@@ V.3 Hardware constraints
In this section, we study how hardware constraints, most notably, the sampling frequency and bandwidth of the arbitrary waveform generator (AWG), can affect the performance of the CZ gate. The hardware limitations are important, because the CZ gates we consider are based on fast-flux control, and it is necessary that current hardware be capable of implementing such fast pulses. Let $F_{s}$ denote the sampling frequency and $bw$ the bandwidth. Examples of state-of-the-art AWGs include QBLOX QCM with $F_{s}=1$ GSa/s and $bw=400$ MHz, Zurich Instrument SHFQC+ with specifications up to $F_{s}=2$ GSa/s and $bw=800$ MHz, QBLOX QCM with $F_{s}=2$ GSa/s and $bw=800$ MHz, Keysight M5300A with baseband sampling frequency $F_{s}=4.8$ GSa/s and $bw=2$ GHz, etc. Details on how we impose the hardware constraints in simulation are summarized in Appendix J.
@@
Index Slepian Chebyshev
$t_{d}$ (ns) $P_{e}$ $t_{d}$ (ns) $P_{e}$
(a) 51.9 $10^{-3.14}$ 51.5 $10^{-3.15}$
(b) 48.5 $10^{-3.91}$ 47.6 $10^{-3.96}$
(c) 47.0 $10^{-4.47}$ 46.0 $10^{-4.55}$
(d) 47.0 $10^{-4.62}$ 46.0 $10^{-4.69}$
Table 2: Best operating points using $\tilde{g}_{\textrm{sl2}}[n]$ and $\tilde{g}_{\textrm{ch2}}[n]$ in Fig. 4. The index corresponds to that in Fig. 4.
We utilize the same comparison example of $\tilde{g}_{\textrm{sl2}}[n]$ and $\tilde{g}_{\textrm{ch2}}[n]$ as in Section V.2.1. Fig. 4 presents the simulation results of the CZ gate using $\tilde{g}_{\textrm{sl2}}[n]$ and $\tilde{g}_{\textrm{ch2}}[n]$ with different practical hardware parameters. The hardware parameters are listed in Table 1. When comparing Fig. 4 against Fig. 3, it becomes evident that there is an overall increase in leakage error, regardless of the control trajectories. The difference between both trajectories also shrinks. As we enhance the hardware parameters, a better resemblance in the performance of the CZ gate to that without hardware constraints is observed. Table 2 shows an aggregate of the best operating points using $\tilde{g}_{\textrm{sl2}}[n]$ and $\tilde{g}_{\textrm{ch2}}[n]$ in Fig. 4 following the same argument as in Section V.2.1. Comparing the best operating points, we argue that the advantage of $\tilde{g}_{\textrm{ch2}}[n]$ over $\tilde{g}_{\textrm{sl2}}[n]$ can be mostly recovered with $F_{s}=5$ GSa/s and $bw=2$ GHz. Thus, the proposal in this paper is readily implementable using off-the-shelf state-of-the-art hardware.
@@ VI Conclusion and outlook
In this work, we formulate the problem of baseband flux control design of an adiabatic CPHASE gate in superconducting circuits as a pulse-design problem, and further as a control-trajectory-design problem. Building upon knowledge from other contexts of pulse-design problems, we propose the Chebyshev-based trajectory as an alternative to the widely used Slepian-based trajectory. We then analytically show the advantage of the Chebyshev-based trajectory by using a two-level system abstraction. Furthermore, we compare the performance of the two types of trajectories by numerically simulating a CZ gate in two capacitively coupled transmon qubits. Our simulation results show that the Chebyshev-based trajectory can be designed to induce lower leakage error by roughly $6\%$ in certain cases, while maintaining similar or even smaller pulse duration. We note that in several cases the Slepian-based trajectory induces lower leakage error due to some abnormal phenomenon not existing in analysis. In addition, we study how practical hardware constraints including sampling frequency and bandwidth can influence the performance of the theoretically derived control pulses and affect the advantage of the Chebyshev-based trajectory over the Slepian-based trajectory in simulation. We find that the advantage can be mostly realized using state-of-the-art hardware.
The design of the flux pulse plays a crucial role in achieving high fidelity and fast speed for baseband flux-based gates in superconducting circuits. The weighted Chebyshev approximation, a versatile technique devised for tailoring pulses based on specific requirements and constraints, emerges as a valuable tool to design pulse shapes. The perspectives in this research could have broader applicability for pulse-design problems of other types of quantum gates in superconducting circuits and other quantum computing hardware platforms.
@@ B.3 Relationship to the Landau-Zener formula
$\log P_{\textrm{LZ}}=\frac{\pi}{4}(\log P_{\textrm{eLZ}}-C)\ .$ (51)
@@ F.1 The Slepian pulses
The Slepian pulses, also known as discrete prolate spheroidal sequences (DPSSs), are a set of orthogonal pulses intended for the problem of maximal concentration in both the time domain and the frequency domain. Heisenberg’s uncertainty principle [52] implies that pulses cannot be confined in both the time domain and the frequency domain. It is then natural to ask: how to optimally concentrate the energy in one domain if the pulse is strictly confined in the other domain. This problem, both in continuous time and in discrete time, was pursued and solved by Slepian, Landau, and Pollack [53, 54, 55, 56, 57, 58]. Here we briefly review the development and analysis of the Slepian pulses for the discrete-time case.
@@ Appendix I Additional simulation results
We present additional simulation comparison results between different pairs of benchmark Slepian-based trajectories $\tilde{g}_{\textrm{sl2}}[n]$ and Chebyshev-based trajectories $\tilde{g}_{\textrm{ch2}}[n]$ in Fig. 12. We follow the simulation and analysis procedure as discussed in Section V. In all these examples, we can see that $\tilde{g}_{\textrm{ch2}}[n]$ pushes the leakage lower in the range of smaller pulse duration while sacrificing a higher leakage error in the range of larger pulse duration. The best operating points with shortest pulse duration are indicated by green squares and purple dots. Note that in Figs. 12(g)-(h), we observe an abnormal small lobe appearing before the common leakage error lobe. Table 5 shows an aggregate of comparisons of best operating points following the same argument as in Section V.2.1.
@@
$t_{d}$ (ns) $P_{e}$ $t_{d}$ (ns) $P_{e}$
(a) 46.8 $10^{-3.09}$ 46.0 $10^{-3.13}$
(b) 46.5 $10^{-3.46}$ 46.0 $10^{-3.49}$
(c) 46.6 $10^{-3.83}$ 46.0 $10^{-3.85}$
(d) 46.7 $10^{-4.18}$ 46.0 $10^{-4.22}$
(e) 47.0 $10^{-4.51}$ 46.5 $10^{-4.52}$
(f) 47.0 $10^{-4.80}$ 46.5 $10^{-4.82}$
(g) 41.5 $10^{-5.75}$ 46.7 $10^{-5.01}$
(h) 42.0 $10^{-7.16}$ 47.0 $10^{-5.15}$
@@
Appendix J Details on simulating hardware limitations

### v2

# Present address: ]Atlantic Quantum, Cambridge, MA 02139

@@ Abstract
Despite progress towards achieving low error rates with superconducting qubits, error-prone two-qubit gates remain a bottleneck for realizing large-scale quantum computers. Therefore, a systematic framework to design high-fidelity gates becomes imperative. One type of two-qubit gate in superconducting qubits is the controlled-phase (CPHASE) gate, which utilizes a conditional interaction between higher energy levels of the qubits controlled by a baseband flux pulse on one of the qubits or a tunable coupler. In this work, we study an adiabatic implementation of CPHASE gates and formulate the design of the control trajectory for the gate as a pulse-design problem. We show in simulation that the Chebyshev-based trajectory can, in certain cases, enable gates with gate infidelity lower by an average of $23.3\%$ when compared to the widely used Slepian-based trajectory.
@@ II.4.2 Constraint on duration
$P_{e}=\frac{\displaystyle\bigg{|}\int\tilde{g}(\tau)e^{-i\Delta\tau}\textrm{d}\tau\bigg{|}^{2}}{4}=\frac{\displaystyle\big{|}G(i\Delta)\big{|}^{2}}{4}\ ,$ (7)
@@ IV Comparison between Chebyshev- and Slepian-based trajectories
We have formulated the CPHASE gate design problem into a pulse design problem and further transformed it into the design of a control trajectory $\tilde{g}[n]$, where the frequency-domain characteristics are emphasized. Special attention is paid to the comparisons between the Chebyshev-based trajectories and the Slepian-based trajectories that will be defined in this section.
@@
To obtain the shortest pulse given a specified allowable leakage error, we need to minimize the frequency cutoff $\omega^{*}$ (see definition in Statement 3) while keeping amplitude of the frequency components beyond $\omega^{*}$ below some constant determined by the allowable leakage error. We propose the Chebyshev-based trajectory $\tilde{g}_{\textrm{ch2}}[n]$ as an alternative to the Slepian-based trajectory $\tilde{g}_{\textrm{sl2}}[n]$. While $\tilde{g}_{\textrm{sl2}}[n]$ can fulfill the desired features because it is originally designed to obtain a maximized energy concentration within some frequency band, it features a decreasing amplitude envelop of sidelobe frequency components, which diminishes its capability to push $\omega^{*}$ even smaller. On the other hand, $\tilde{g}_{\textrm{ch2}}[n]$, by design, features a flat amplitude envelop of sidelobe frequency components, which allows $\omega^{*}$ to be even smaller. The argument is that we can allow higher sidelobe amplitudes in larger frequency components as long as they stay below a specified threshold, and therefore we can in turn decrease the concentration in smaller frequency components. In this way, $\tilde{g}_{\textrm{ch2}}[n]$ can be designed to obtain a smaller $\omega^{*}$ while maintaining a sidelobe amplitude below some threshold. Here, $\omega^{*}$ denotes the smallest attainable frequency under certain constraints given some leakage threshold, as we specify in Statement 3. This eventually leads to the fact that the Chebyshev-based trajectory can be designed to be shorter than the Slepian-based trajectory.
@@ V.1 Setting and procedure
In our simulation, we choose the parameters $\omega_{2}/2\pi=4.7$ GHz, $\alpha_{1}/2\pi=\alpha_{2}/2\pi=-300$ MHz, and $g/2\pi=10\sqrt{2}\approx 14.142$ MHz. We set $\omega_{1}/2\pi=5.8$ GHz initially and tune $\omega_{1}$ to perform a CZ gate. These parameters form a typical parameter set for transmon qubits and the operation of a CZ gate. In order to perform a CZ gate, we detune $\omega_{1}$ so that $\omega_{1}+\omega_{2}\approx 2\omega_{1}+\alpha_{1}$. If we move exactly to the degeneracy point of diabatic states |$\ket{11}$ and |$\ket{20}$, then we have $\omega_{1}/2\pi=5.0$ GHz. We vary $\omega_{1}/2\pi$ from $5.8$ GHz to approximately $5.0$ GHz depending on the amplitude of the control pulse, and then back to $5.8$ GHz.
@@ V.2.1 A comparison example
Potential operating points are considered to be the insensitive points to pulse duration, i.e., the points of the leakage error lobes as shown by the green squares and purple dots in Fig. 3. In addition, if there is a slight deviation in pulse duration at those points, we would be almost surely have a lower leakage error regardless of the direction of the deviation. We determine the best operating point to be the one with the smallest pulse duration, indicated by the red circle in Fig. 3. Since $\tilde{g}_{\textrm{ch2}}[n]$ generally pushes the leakage error lower in the range of relatively smaller pulse duration, we are able to achieve an operating point with lower leakage error while also maintaining a similar or even smaller pulse duration compared to $\tilde{g}_{\textrm{sl2}}[n]$. The operating points for $\tilde{g}_{\textrm{sl2}}[n]$ and $\tilde{g}_{\textrm{ch2}}[n]$ are $t_{d}=47.0$ ns with $P_{e}=10^{-4.66}$ and $t_{d}=46.1$ ns with $P_{e}=10^{-4.72}$, respectively. We further compute the average gate fidelity as described in Appendix J, and find that the infidelity $1-F_{g}=5.5\times 10^{-6}$ for $\tilde{g}_{\textrm{sl2}}[n]$ and $1-F_{g}=6.8\times 10^{-6}$ for $\tilde{g}_{\textrm{ch2}}[n]$ respectively. In this case, the average gate infidelity of the Slepian-based trajectory is lower than that of the Chebyshev-based trajectory although the leakage error of the former is greater than that of the latter. This is mainly because the unwanted population exchange between |$\ket{01}$ and |$\ket{10}$ within the computational subspace contributes to the average gate infidelity.
@@ V.2.2 An aggregate of comparison examples
We observe similar results in terms of leakage error to that in Section V.2.1 for a majority of the comparison examples. We find that $\tilde{g}_{\textrm{ch2}}[n]$ can be designed to achieve a lower leakage error than its counterpart $\tilde{g}_{\textrm{sl2}}[n]$ by roughly $6.0\%$ in certain cases, while also maintaining smaller pulse duration by an average of $0.6$ ns. In several unusual cases, where there appears an unusually small leakage lobe before the first main leakage lobe, $\tilde{g}_{\textrm{sl2}}[n]$ in fact induces lower leakage error with smaller pulse duration. This indicates that simulation verification is important once a control trajectory is designed because there can be discrepancies between analysis and simulation due to some approximation. When it comes to gate infidelity, for most of the comparison examples ((a)-(g)), we find that the Chebyshev-based trajectories have lower gate infidelity than the Slepian-based trajectories. In example (h), the Slepian-based trajectory has lower gate infidelity because the leakage error is significantly smaller, due to the fact that there appears a significantly small leakage lobe before the first main leakage lobe. On average, the Chebyshev-based trajectories achieve a $23.3\%$ reduction in gate infidelity. This can also be interpretated as an increase in gate fidelity by $0.0009\%$. We note that this small number in terms of increase in gate fidelity results from the fact that the Slepian-based trajectories have already reached a relatively high fidelity (close to $1$). In other words, the percentage is almost identical to the absolute increase in gate fidelity. To be more specific, given that a Slepian-based trajectory achieves $x\%$ gate fidelity, the corresponding Chebyshev-based trajectory can on average achieves approximately $(x+0.0009)\%$ gate fidelity. We note that this improvement can become more prominent when the desired gate fidelity is higher (e.g., $x>99.999$). This demonstrates that the Chebyshev-based trajectories can potentially enhance overall gate performance even when accounting for more than leakage error.
To determine scenarios where Chebyshev-based trajectories might significantly outperform other standard approaches, we find that in the region where the Slepian-based trajectory achieves a gate infidelity of $\sim 10^{-3}$ to $\sim 10^{-5}$, the Chebyshev-based trajectory appears to consistently decrease the infidelity by roughly $40.7\%$. However, in the infidelity range below $\sim 5\times 10^{-6}$, we see mixed behavior and it becomes hard to conclude for sure which one is better. This indicates that the Chebyshev-based trajectory might outperform other standard approaches in relatively moderate infidelity ranges, but may not be as effective if the desired infidelity is very low.
We note that there can be different operating points, and in our comparison, we further pick the one with the shortest pulse duration to be the best operating point because this aligns with our definition of a desired pulse in the first place. One could potentially pick any of these operating points as their operating points if a longer pulse duration is acceptable, and the comparison results will change accordingly. However, our approach of best operating point aims to avoid any bias in the evaluation.
@@ V.3 Hardware constraints
In this section, we study how hardware constraints, most notably, the sampling frequency and bandwidth of the arbitrary waveform generator (AWG), can affect the performance of the CZ gate. The hardware limitations are important, because the CZ gates we consider are based on fast-flux control, and it is necessary that current hardware be capable of implementing such fast pulses. Let $F_{s}$ denote the sampling frequency and $bw$ the bandwidth. Examples of state-of-the-art AWGs include QBLOX QCM with $F_{s}=1$ GSa/s and $bw=400$ MHz, Zurich Instrument SHFQC+ with specifications up to $F_{s}=2$ GSa/s and $bw=800$ MHz, QBLOX QCM with $F_{s}=2$ GSa/s and $bw=800$ MHz, Keysight M5300A with baseband sampling frequency $F_{s}=4.8$ GSa/s and $bw=2$ GHz, etc. Details on how we impose the hardware constraints in simulation are summarized in Appendix K.
@@
We utilize the same comparison example of $\tilde{g}_{\textrm{sl2}}[n]$ and $\tilde{g}_{\textrm{ch2}}[n]$ as in Section V.2.1. Fig. 4 presents the simulation results of the CZ gate using $\tilde{g}_{\textrm{sl2}}[n]$ and $\tilde{g}_{\textrm{ch2}}[n]$ with different practical hardware parameters. The hardware parameters are listed in Table 1. When comparing Fig. 4 against Fig. 3, it becomes evident that there is an overall increase in leakage error, regardless of the control trajectories. The difference between both trajectories also shrinks. As we enhance the hardware parameters, a better resemblance in the performance of the CZ gate to that without hardware constraints is observed. In Appendix I, we show an aggregate of the best operating points using $\tilde{g}_{\textrm{sl2}}[n]$ and $\tilde{g}_{\textrm{ch2}}[n]$ in Fig. 4 following the same argument as in Section V.2.1. Comparing the best operating points, we argue that the advantage of $\tilde{g}_{\textrm{ch2}}[n]$ over $\tilde{g}_{\textrm{sl2}}[n]$ can be mostly recovered with $F_{s}=5$ GSa/s and $bw=2$ GHz. Thus, the proposal in this paper is readily implementable using off-the-shelf state-of-the-art hardware.
@@ VI Conclusion and outlook
In this work, we formulate the problem of baseband flux control design of an adiabatic CPHASE gate in superconducting circuits as a pulse-design problem, and further as a control-trajectory-design problem. Building upon knowledge from other contexts of pulse-design problems, we propose the Chebyshev-based trajectory as an alternative to the widely used Slepian-based trajectory. We then analytically show the advantage of the Chebyshev-based trajectory by using a two-level system abstraction. Furthermore, we compare the performance of the two types of trajectories by numerically simulating a CZ gate in two capacitively coupled transmon qubits. Our simulation results show that the Chebyshev-based trajectory can be designed to induce lower leakage error by roughly $6.0\%$ in certain cases, while maintaining similar or even smaller pulse duration. We note that in several cases the Slepian-based trajectory induces lower leakage error due to some unusual phenomenon not existing in analysis. On average, we find that the Chebyshev-based trajectory results in $23.3\%$ lower gate infidelity than its Slepian counterpart. From the perspective of quantum error correction, it is the reduction in gate infidelity—rather than the absolute increase in gate fidelity—that more directly impacts the efficacy of error correction protocols. The $23.3\%$ average reduction in gate infidelity of the Chebyshev-based trajectories is considerable, as physical error rates are exponentially leveraged in the suppression of logical error rates [48]. This, in turn, can help reduce the required code distance for achieving a target logical error rate. In addition, we study how practical hardware constraints including sampling frequency and bandwidth can influence the performance of the theoretically derived control pulses and affect the advantage of the Chebyshev-based trajectory over the Slepian-based trajectory in simulation. We find that the advantage can be mostly realized using state-of-the-art hardware.
The design of the flux pulse plays a crucial role in achieving high fidelity and fast speed for baseband flux-based gates in superconducting circuits. The weighted Chebyshev approximation, a versatile technique devised for tailoring pulses based on specific requirements and constraints, emerges as a valuable tool to design pulse shapes. Future work could explore extending this framework to other types of quantum gates, integrating these pulse design techniques into closed-loop calibration and optimal control pipelines, and studying the robustness of the gate performance against control noise in experimental deployment.
@@ Acknowledgements.
Data Availability
The data that support the findings of this article are openly available [49].
@@ B.3 Relationship to the Landau-Zener formula
$\log P_{\textrm{LZ}}=\frac{\pi}{4}(\log P_{\textrm{eLZ}}-C)\ ,$ (51)
@@ F.1 The Slepian pulses
The Slepian pulses, also known as discrete prolate spheroidal sequences (DPSSs), are a set of orthogonal pulses intended for the problem of maximal concentration in both the time domain and the frequency domain. Heisenberg’s uncertainty principle implies that pulses cannot be confined in both the time domain and the frequency domain. It is then natural to ask: how to optimally concentrate the energy in one domain if the pulse is strictly confined in the other domain. This problem, both in continuous time and in discrete time, was pursued and solved by Slepian, Landau, and Pollack [54, 55, 56, 57, 58, 59]. Here we briefly review the development and analysis of the Slepian pulses for the discrete-time case.
@@ Appendix I Additional simulation results
We present additional simulation comparison results between different pairs of benchmark Slepian-based trajectories $\tilde{g}_{\textrm{sl2}}[n]$ and Chebyshev-based trajectories $\tilde{g}_{\textrm{ch2}}[n]$ in Fig. 12. We follow the simulation and analysis procedure as discussed in Section V. In all these examples, we can see that $\tilde{g}_{\textrm{ch2}}[n]$ pushes the leakage lower in the range of smaller pulse duration while sacrificing a higher leakage error in the range of larger pulse duration. The best operating points with shortest pulse duration are indicated by green squares and purple dots. Note that in Figs. 12(g)-(h), we observe an unusually small lobe appearing before the first main leakage error lobe. Table 4 shows an aggregate of comparisons of best operating points following the same argument as in Section V.2.1.
We also present additional simulation comparison results under certain hardware constraints. Table 5 shows an aggregation of the best operating points using $\tilde{g}_{\textrm{sl2}}[n]$ and $\tilde{g}_{\textrm{ch2}}[n]$ in Fig. 4 following the same argument as in Section V.2.1.
@@
$t_{d}$ (ns) $\log(P_{e})$ $1-F_{g}$ $t_{d}$ (ns) $\log(P_{e})$ $1-F_{g}$
(a) 46.8 $-3.09$ $1.8\times 10^{-4}$ 46.0 $-3.13$ $1.5\times 10^{-4}$
(b) 46.5 $-3.46$ $8.4\times 10^{-5}$ 46.0 $-3.49$ $6.8\times 10^{-5}$
(c) 46.6 $-3.83$ $3.7\times 10^{-5}$ 46.0 $-3.85$ $1.8\times 10^{-5}$
(d) 46.7 $-4.18$ $1.7\times 10^{-5}$ 46.0 $-4.22$ $9.5\times 10^{-6}$
(e) 47.0 $-4.51$ $7.3\times 10^{-6}$ 46.5 $-4.52$ $4.0\times 10^{-6}$
(f) 47.0 $-4.80$ $4.2\times 10^{-6}$ 46.5 $-4.82$ $2.9\times 10^{-6}$
(g) 41.5 $-5.75$ $6.8\times 10^{-6}$ 46.7 $-5.01$ $2.2\times 10^{-6}$
(h) 42.0 $-7.16$ $1.2\times 10^{-6}$ 47.0 $-5.15$ $1.7\times 10^{-6}$
@@
Index Slepian Chebyshev
$t_{d}$ (ns) $log(P_{e})$ $1-F_{g}$ $t_{d}$ (ns) $log(P_{e})$ $1-F_{g}$
(a) 51.9 $-3.14$ $1.5\times 10^{-4}$ 51.5 $-3.15$ $1.4\times 10^{-4}$
(b) 48.5 $-3.91$ $2.5\times 10^{-5}$ 47.6 $-3.96$ $1.8\times 10^{-5}$
(c) 47.0 $-4.47$ $7.9\times 10^{-6}$ 46.0 $-4.55$ $9.5\times 10^{-6}$
(d) 47.0 $-4.62$ $5.9\times 10^{-6}$ 46.0 $-4.69$ $7.5\times 10^{-6}$
Table 5: Best operating points using $\tilde{g}_{\textrm{sl2}}[n]$ and $\tilde{g}_{\textrm{ch2}}[n]$ under certain hardware constraints in Fig. 4. The index corresponds to that in Fig. 4.
Appendix J Average gate fidelity
[…]
$F_{g}=\frac{d_{1}F_{p}+1-L_{1}}{d_{1}+1}.$ (88)
We define gate infidelity to be $1-F_{g}$.
Appendix K Details on simulating hardware limitations